In a triangle, the second angle is twice the first and the third angle is 20 drgees less than the first. Find the measure of all three angles.
Thank You!

Answer by venugopalramana(3286)

In a triangle, the second angle is twice the first and the third angle is 20 drgees less than the first. Find the measure of all three angles.
LET FIRST ANGLE =A
SECOND ANGLE = 2*A
THIRD ANGLE = A-20
SUM OF ALL 3 ANGLES =A+2A+A-20=180
4A=180+20=200
A=200/4=50
SO THE 3 ANGLES ARE A=50,2*50=100,50-20=30.